OpenStudy (anonymous):

What is the sum of five over six plus one over three in simplest form? one and one over six one and one over three seven over eight two over three

OpenStudy (anonymous):

@*Shadow* @confluxepic

OpenStudy (confluxepic):

\[\frac{ 5 }{ 6 }+\frac{ 1 }{ 3 }\]

OpenStudy (confluxepic):

\[\frac{ 5 }{ 6 }+\frac{ 2 }{ 6 }\]

OpenStudy (confluxepic):

\[\frac{ 7 }{ 6 }\]

OpenStudy (confluxepic):

\[1\frac{ 1 }{ 6}\]

OpenStudy (confluxepic):

One and one over six.